at certain pressure and at 17 C temperature the fixed amount of a gas has volume of 0.73L . without change in pressure at what temperature the volume of gas will be 0.85L
1
Expert's answer
2013-01-25T06:47:56-0500
You can find temperature by using next law:

(P1V1)/T1 = (P2V2)/T2

P1 = const ( without change in pressure )
V1 = 0.73 L
T1 = 17 C
V2 = 0.85 L
T2 - ?

If there is no change in pressure
V1/T1 = V2/T2

T2 = 0.85 / (0.73/17)

T2 = 19.81 C